OpenStudy (desiraem):

how would you solve this problem? i'm trying to learn for my PERT test. :) 9 ∙ 9 - 8 ∙ 3 + (-30)

OpenStudy (anonymous):

I would do pemdas (Parenthesis, Exponents, Multiplication, Division, Addition and Subtraction). Except, cross out the ones that aren't there. So I would MAS, which is Multiplication, Addition and subtraction. So for example, 9*9=81 81-8= 73 73*3= 219 219+(-30)= 189 Hopefully this helped.

OpenStudy (desiraem):

i tried the pemdas but, the answer wasn't on my answer choices so i'm stuck lol :( -27 87 75 27

OpenStudy (anonymous):

I would e-mail a math teacher than and have them go over some of the problems with you and give you some example ones to figure it out. But that is a trick question though. :\

OpenStudy (desiraem):

lol it's super fine! thank you for trying :)